# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

### v2

@@ Abstract
Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability by making the agent’s decisions understandable and verifiable. Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and may fail to generalize to unseen environments. We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization. iTRACE first extracts semantically meaningful features, then maps them to policies via interpretable trees. To address the impracticality of running VLMs in RL loops, we distill their outputs into a lightweight model. By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE eliminates the need for human annotation traditionally required by interpretable models, while also addressing the limitations of VLMs alone, such as their lack of grounding in action spaces and inability to directly optimize policies. iTRACE outperforms MLP baselines that use the same interpretable features and matches the performance of CNN-based policies, producing verifiable, semantically interpretable, and human-aligned behaviors without requiring human annotation.
1 Introduction
@@
Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [28]. Ensuring that these learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions. While progress has been made in developing interpretable AI models, achieving semantic interpretability in RL remains particularly challenging  [22].
Semantic interpretability refers to the extent to which a model’s features, decisions, or internal representations can be understood and explained in terms of human-understandable concepts. Achieving this requires two essential components: (1) the inputs to the decision-making model must be intelligible to facilitate downstream interpretation, and (2) the decision-making process must be built with an interpretable model that allows humans to understand and explain its rationale  [9]. In essence, achieving semantic interpretability involves mapping high-dimensional input data to human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models.
@@
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions. Decision trees have been used to represent policies in a structured, human-readable format [30, 26, 18]. Rule-based methods offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 12]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
Existing approaches have attempted to bridge the gap between raw image inputs and interpretable features [2] to facilitate model understanding but fall short in semantic interpretability. Many approaches either produce mathematical or visual representations that do not correspond to human-understandable concepts [15, 10] or rely on predefined feature spaces that aren’t generalize across different environments, using CNNs [8] or ViTs [23] to establish mappings between inputs and semantic features. As a result, human intervention has traditionally played a key role, either by manually defining semantically meaningful feature spaces [40] or by labeling machine-generated features to associate them with semantically meaningful concepts [29].
The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges. These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [4, 38, 6], have seen widespread use in various tasks  [38, 39]. In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
@@
We propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments. By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts. Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction.
2 Related work
2.1 Interpretable reinforcement learning
@@
Figure 1: Overview of iTRACE, our proposed framework. A. Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation. (2) The VLM generates code to extract features from a segmentation of a single image frame and determines whether certain key features require multiple frames. If so, it extends the single-frame extraction code to handle multi-frame inputs. The segmentation is performed using the Segment Anything Model (SAM). (3) An exploration policy is used to generate a dataset of feature–raw image frame pairs by applying SAM and the feature extraction code. To avoid the computational overhead of running SAM at every reinforcement learning step, we train an Extractor Module as an efficient substitute. B. Tree-based RL Training leverages the pretrained Extractor Module to obtain interpretable features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
Interpretable RL is inherently challenging compared to standard supervised learning, as it needs to account for the long-term consequences of actions, partial observability, and sparse feedback during exploration [22]. Post-hoc methods like Jacobian-based, perturbation-based, and attribution-based saliency maps  [25, 10, 34] offer limited interpretability, as they lack transparency in both inputs and decision-making [9]. Approaches to achieve intelligible decision-making include rule lists and decision trees [9], among which decision trees are arguably the easiest to verify  [27]. Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) support gradient-based optimization of interpretable models [27], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, balancing performance with interpretability [19]. Similarly, POETREE (Policy Extraction through Decision Trees) [17] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements. Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs
2.2 Feature extraction for reinforcement learning
Nonparametric methods (e.g., manifold and spectral learning) and nonlinear parametric methods (e.g., deep learning) are commonly used in RL for feature extraction but typically do not produce interpretable features [15]. Other work focuses on the extraction of visual features that improve performance but do not guarantee human-interpretable concepts. Yet, few works attempt to construct semantically meaningful feature spaces without human supervision for RL training.  [35] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding objects in the environment. However, they report that the extracted features could be noisy and inconsistent  [35]. Our framework differs from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
2.3 VLMs for reinforcement learning
VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation extraction [21, 33, 6]. One prominent approach involves using VLMs to generate reward functions.  [21] use VLMs as zero-shot reward models to specify tasks via natural language, while  [32] leverages VLMs to produce dense reward functions through code generation. Additionally,  [33, 31] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL. Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.  [6] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies. However, these representations are not guaranteed to be accurate or semantically meaningful. To our knowledge, this is the first approach that uses VLMs to automatically construct semantically interpretable feature spaces for RL.
3 Methodology
We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment. iTRACE leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight vision model (which we refer to as the Extractor Module) as a computationally efficient substitute. Subsequently, utilizing the Extractor Module to replicate VLM feature extraction, iTRACE performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
3.1 Feature extraction
The feature extraction process of iTRACE involves a three-step pipeline, as shown in Fig. 1. In Step 1, we query the VLM to generate key features that contribute to the decision in the environment. In Step 2, we query the VLM to generate code that extracts the identified features, either from a single image or a list of images depending on whether the features demand temporal context. Both raw observations and segmentation masks are provided to the VLM in this process. In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor Module to increase program speed in RL training. The methodology is formalized below.
Step 1: Querying the VLM for Semantic Features.
@@ 3.1 Feature extraction
To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making. In this step, we provide the VLM with:
  • $\mathcal{D}$: Basic domain knowledge about the environment.
@@
For example, in the Skiing environment, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
@@
We instruct the VLM to provide a "requires_temporal_context" field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$. While $\mathcal{F}_{m}$ may be unnecessary in some environments, it is essential in others. Returning to the Skiing example, the skier must actively tilt left and right to maintain speed. Without knowledge of the skier’s motion direction, it is ambiguous which way the skier should tilt. We allow the VLM to determine whether temporal context is required for effective decision-making in a given environment. A complete list of extracted features is provided in Appendix A.2.2.
Step 2: Querying the VLM for Feature Extraction Code.
@@
In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$ from one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. Recognizing that hallucination is a well-documented challenge in generative VLM models [16] (see Appendix. A.5), we use segmentation to provide accurate spatial information and ground the generated extraction code in accurate visual data. This strategy aligns with common practices in the field [20, 37]. The visual segmentation $\mathcal{S}$ is computed using Segment Anything [13], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. We use the ViT-H SAM model with its default parameters and no modifications to ensure the generalizability of our approach.
@@
Example Segmentations Labeling: While segmentation improves spatial grounding, its interpretation by the VLM remains prone to hallucination. To ensure the generated code generalizes across diverse environment frames, the VLM must be exposed to a representative set of example images. However, the raw JSON format of segmentation can lead to confusion between segments across different images. To address this, we first prompt the VLM to label a small set of segmentations associated with example frames, processing one image at a time. An example of this labeling process is shown in Fig. 9(b), where all segment labels are generated by the VLM based on raw segmentation masks.
@@
Following this, depending on whether the feature set requires temporal context, the VLM is instructed to take one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. The labeled segmentations serve as reference examples, helping the VLM infer the most likely attribute values of relevant objects and formulate logic to filter out noisy segments. The code generation process involves two steps:
Single Image Extraction: For features in $\mathcal{F}_{s}$, the VLM generates code $\mathcal{G}_{s}$ to extract their values from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$. To enhance the robustness of the generated code, the VLM is provided with multiple example images and their labeled segmentation during this stage to ensure it has been exposed to most possible scenarios. See Algorithm 1. Also see Fig. 9(c) for an example output on segmentation masks in the Skiing environment.
Multiple Image Extraction: For features in $\mathcal{F}_{m}$, if any, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$. Multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$ are provided to the VLM during generation. See Algorithm 2.
The final feature extraction code $\mathcal{G}$ is defined in Eq. 2:
$\mathcal{G}=\begin{cases}\mathcal{G}_{s},&\text{if }\mathcal{F}_{m}=\emptyset;\\ \mathcal{G}_{m}(\mathcal{G}_{s}),&\text{if }\mathcal{F}_{m}\neq\emptyset.\end{cases}$ (2)
Algorithm 1 Generating Extraction Code: Single Image
1:  Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, VLM $\mathcal{M}$
2:  Initialize $\mathcal{G}_{s}=null$
3:  $examples\leftarrow\emptyset$
4:  for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:     $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:     Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:  end for
8:  $\mathcal{G}_{s}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples)$
Algorithm 2 Generating Extraction Code: Multiple Images
1:  Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, $\mathcal{G}_{s}$, VLM $\mathcal{M}$
2:  Initialize $\mathcal{G}_{m}=null$
3:  $examples\leftarrow\emptyset$
4:  for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:     $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:     Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:  end for
8:  $\mathcal{F}_{m}\leftarrow{\mathcal{M}_{\text{JudgeMultiFrame}}}(\mathcal{F})$
9:  if $\mathcal{F}_{m}\neq\emptyset$ then
10:     $\mathcal{G}_{m}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples,\mathcal{G}_{s})$
11:  end if
12:  Return: $\mathcal{G}_{m}$
@@
Figure 2: An example of how raw image observations are processed in our pipeline. (a) Raw image observations are segmented using Segment Anything. (b) A VLM assigns semantic labels to segments in a small set of example images. (c) A VLM then generates code that extracts target features from segmentation masks, applicable to all images. Step (b) is only performed on the example set to guide code generation.
Step 3: Dataset Generation and Training the Extractor Module.
@@
In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$.
@@
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$. The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training. In our experiments, we use a CNN-based Extractor Module (architecture included in Appendix A.3) due to its low computational overhead in simulated environments. However, the framework is compatible with any vision model, including fine-tuned ViT transformers or other task-specific architectures based on deployment needs.
@@
If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$ The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude.
3.2 Tree-based RL training
After obtaining the Extractor Module, we train an interpretable tree-based model with RL to connect the semantic features with action outputs. During RL training, our framework utilizes the pre-trained Extractor Module to extract semantic features from image observations, as shown in Fig. 1. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step. Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
3.2.1 Interpretable control tree

### v3

@@ Abstract
Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability of decision-making. Achieving semantic interpretability in reinforcement learning requires (1) a feature space composed of human-understandable concepts and (2) a policy that is interpretable and verifiable. However, constructing such a feature space has traditionally relied on manual human specification, which often fails to generalize to unseen environments. Moreover, even when interpretable features are available, most reinforcement learning algorithms employ black-box models as policies, thereby hindering transparency. We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and train a interpretable tree-based model via RL. To address the impracticality of running VLMs in RL loops, we distill their outputs into a lightweight model. By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE loosens the reliance the need for human annotation that is traditionally required by interpretable models. In addition, it addresses key limitations of VLMs alone, such as their lack of grounding in action spaces and their inability to directly optimize policies. We evaluate iTRACE across three domains: Atari games, grid-world navigation, and driving. The results show that iTRACE outperforms other interpretable policy baselines and matches the performance of black-box policies on the same interpretable feature space.
I Introduction
@@
Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [1]. Ensuring the learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions. While progress has been made in interpretable AI, achieving semantic interpretability in RL remains particularly challenging [2]. It requires (1) constructing a feature space composed of human-understandable concepts, a process that has traditionally relied on human annotation and lacks generalizability. (2) In addition, it requires learning policies that are both high-performing and transparent [3].
@@
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions [3]. Decision trees have been used to represent policies in a structured, human-readable format [4, 5]. Rule-based methods offer an alternative by encoding policies in a form that can be explicitly followed and verified [6, 7]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
Figure 1: High-level overview of iTRACE.
In this paper, we tackle the challenge of mapping high-dimensional observations into human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models. Prior work has attempted to bridge the gap between raw image inputs and interpretable features [8], but existing methods fall short of true semantic interpretability. Some approaches generate mathematical or visual representations that do not correspond to human-understandable concepts [9, 10]. Others rely on predefined feature spaces that fail to generalize across environments [11, 12]. As a result, human involvement remains central: researchers either manually define semantic features [13] or label machine-generated ones [14]. However, this reliance on human annotation is not automated and incurs significant time costs, limiting scalability to new domains.
The emergence of Vision-Language Models (VLMs) has opened new possibilities for addressing these challenges, particularly the difficulty of building human-aligned feature extractor. VLMs, known for their capabilities in visual understanding, language reasoning, and decision-making  [15], have seen widely used in various tasks  [15, 16].
@@
In this work, we propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE). iTRACE leverages VLMs to extract semantic features from raw pixel inputs and trains Interpretable Control Trees to provide the rationale behind RL decisions in terms of human-understandable concepts. Our framework (i) achieves performance competitive with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both single-frame and multi-frame inputs, enabling the capture of temporal information, (iii) generalizes to unseen environments without human involvement, and (iv) ensures computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction. To the best of our knowledge, iTRACE is the first automated framework to achieve semantic interpretability in RL while generalizing effectively to new environments without human-in-the-loop.
II Related work
II-A Interpretable reinforcement learning
@@
Figure 2: Overview of iTRACE, our proposed framework. A. Feature Extraction follows three steps: (1) the VLM identifies key semantic features from domain knowledge and observations; (2) it generates code to extract features from SAM-based segmentations, extending to multi-frame inputs if needed; and (3) an exploration policy generates feature–image pairs, which are used to train a lightweight Extractor Module to replace costly SAM queries during RL. B. Tree-based RL Training then leverages the Extractor Module to obtain interpretable features, which serve as inputs to an Interpretable Control Tree for policy learning.
Interpretable RL remains an open challenge, as current methods struggle to combine human-understandable policies with high performance, to determine when such policies are feasible, and to maintain transparency in high-dimensional state spaces [2]. Following Rudin’s definition, we view Interpretable RL as learning constrained models whose reasoning humans can directly understand, distinct from post-hoc explainability tools such as SHAP [17], LIME [18], or saliency maps [19, 20], which are often misleading and susceptible to hallucinations.
Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs. Among inherently interpretable models, tree-based methods stand out for producing explicit human-readable rules [21], handling categorical and nonlinear interactions [22], supporting sparsity constraints [23], and achieving accuracy competitive with black-box models when properly optimized [24, 25, 26]. Recent advances extend these strengths: Differentiable Decision Trees (DDTs) enable gradient-based optimization of interpretable models [27]; Interpretable Continuous Control Trees (ICCTs) extend this approach to continuous control tasks, balancing transparency and performance [28]; and POETREE introduces dynamically growing probabilistic trees that evolve in complexity during training [29].
II-B Feature extraction for reinforcement learning
Nonparametric methods (e.g., manifold and spectral learning) and nonlinear parametric methods (e.g., deep learning) are commonly used in RL for feature extraction but typically do not produce interpretable features [9]. Other approaches focuses on identifying visual representations that boost performance without guaranteeing human-meaningful concepts: SSINet highlights task-relevant regions with attention masks [19], while perturbation-based saliency maps identify influential pixels [10]. Yet, few works construct semantically meaningful feature spaces without human supervision for RL training. For example, Wu et al. [30] automatically parse features from Atari manuals, but their extracted features remain noisy and still rely on oracle human-defined information. In contrast, our framework automatically defines and extracts features without human-in-the-loop assumptions.
III Interpretable Feature Extraction
In this section, we introduce the interpretable feature extraction component of the iTRACE framework, shown in the first block of Fig. 2. This component leverages the visual comprehension and reasoning capabilities of VLMs to automate feature extraction through a three-step pipeline: (1) zero-shot prompting is used to identify semantic features that deterministically influence decision-making (line 3 in Algorithm 1); (2) feature extraction code is generated to capture both static and temporal information (line 4 in Algorithm 1). Both raw observations and segmentation masks are provided to the VLM in this process; and (3) a dataset is collected using the VLM-generated code to train a lightweight vision model (the Extractor Module) as a computationally efficient substitute (line 5 and line 6 in Algorithm 1). Step 3 is crucial because VLM-generated code depends on image segmentation, which is prohibitively slow and non-batchable when performed at every RL step, resulting in impractical training times.
III-A Deduce features
@@
To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making (line 3 in Algorithm 2). In this step, we provide the VLM with:
  • $\mathcal{D}$: Domain knowledge limited to basic information, obtainable via foundation model queries or agentic exploration (Sec. VI).
@@
For example, in the Skiing environment, as deduced by the VLM, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
@@
We instruct the VLM to give a "temporal_context" boolean field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$. While $\mathcal{F}_{m}$ may be unnecessary in some environments, it is essential in others. Returning to the Skiing example, the skier must actively tilt left and right to maintain speed. Without knowledge of the skier’s motion direction, it is ambiguous which way the skier should tilt. We allow the VLM to determine whether temporal context is required for effective decision-making in a given environment.
III-B Generate feature extraction code
@@
In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$ from one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. Recognizing that hallucination is a well-documented challenge in generative VLM models [31], we use segmentation to provide accurate spatial information and ground the generated extraction code in visual data. This strategy aligns with common practices in the field [32, 33]. The visual segmentation $\mathcal{S}$ is computed using Segment Anything Model (SAM) [34], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. We use the ViT-H SAM model with its default parameters and no modifications to ensure the generalizability of our approach.
@@
Example Segmentations Labeling: While segmentation improves spatial grounding, its interpretation by the VLM remains prone to hallucination. To ensure the generated code generalizes across diverse environments, the VLM must be exposed to a representative set of example images. However, providing multiple segmentation JSONs at once can cause the VLM to conflate objects across images, consistent with known failures of VLMs to maintain object identity across contexts [35]. To address this, we first prompt the VLM to label a small set of segmentations associated with example frames, processing one image at a time. An example of this labeling process is shown in Fig. 4(b), where all segment labels are generated by the VLM based on raw segmentation masks.
@@
Following this, depending on whether the feature set requires temporal context, the VLM is instructed to take one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. The labeled segmentations serve as reference examples, helping the VLM infer the most likely attribute values of relevant objects and formulate logic to filter out noisy segments. The code generation process is described in line 4-9 in Algorithm 2), where the VLM first generates code to extract values from a single image–segmentation pair $(\mathcal{I},\mathcal{S})$, and then generates code $\mathcal{G}_{m}$, extending $\mathcal{G}_{s}$ to handle sequences of image-segmentation pair $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$.
Algorithm 1 iTRACE Overall Framework
1:Input: Environment $\mathcal{E}$, VLM $\mathcal{M}$, RL algorithm (PPO), horizon $H$
2: // Call perception stack
[…]
7: // Call RL stack
[…]
9:Return: trained ICT policy $\pi_{\theta}$
Algorithm 2 Perception Stack: Automated Feature Extraction
1:Input: Environment $\mathcal{E}$, VLM $\mathcal{M}$, Domain $\mathcal{D}$
2: Sample example images $\{I_{i}\}_{i=1}^{n}$ from $\mathcal{E}$ // Step 1: Deduce features
[…]
5:if $F_{m}\neq\emptyset$ then
6:  $G_{m}\leftarrow\mathcal{M}_{\text{Generate}}(F_{m},G_{s})$
7:  $G\leftarrow(G_{s},G_{m})$
8:else
9:  $G\leftarrow G_{s}$
10:end if// Step 3: Train extractor module
11: Collect dataset $\mathcal{T}=\{(I_{t},G(I_{t}))\}$ via exploration
12: Train lightweight extractor $\phi$ on $\mathcal{T}$
13:Return: Extractor module $\phi$
Algorithm 3 RL Stack: Training Interpretable Control Tree
1:Input: Extractor module $\phi$, environment $\mathcal{E}$, horizon $H$, PPO parameters
2: Initialize ICT policy $\pi_{\theta}$
3:for episode $1,\dots,N$ do
4:  Reset $\mathcal{E}$, observe raw frame $I_{0}$
5:  for t = 1 to H do
6:   $f_{t}\leftarrow\phi(I_{t})$ // semantic features
7:   $a_{t}\sim\pi_{\theta}(f_{t})$
8:   Apply $a_{t}$ in $\mathcal{E}$, receive $I_{t+1},r_{t}$
9:  end for
10:  Update $\pi_{\theta}$ via PPO on collected $(f_{t},a_{t},r_{t})$
11:end for
12:Return: trained ICT policy $\pi_{\theta}$
@@
Figure 3: An example of how raw image observations are processed in our pipeline. (a) Raw image observations are segmented using SAM. (b) A VLM assigns semantic labels to segments in a small set of example images. (c) A VLM then generates code that extracts target features from segmentation masks, applicable to all images. Step (b) is only performed on the example set to guide code generation.
III-C Dataset generation and training the extractor module.
In this step, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently. This non-batchability severely impacts downstream RL training.
@@
We execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images, $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$ (line 11 in Algorithm 2).
@@
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$ (line 12 in Algorithm 2). The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training. In our experiments, we use a CNN-based Extractor Module due to its low computational overhead in simulated environments. While the framework is compatible with any lightweight vision model, including fine-tuned ViTs or task-specific architectures, CNNs provide an effective and widely validated baseline that balances accuracy, efficiency, and robustness. The execution time comparison between VLM-generated feature extraction code and our CNN-based Extractor Module is demonstrated in Table I.
TABLE I: Runtime speedup on Skiing. Similar results achieved on other domains.
Batch VLM Code CNN Speedup
[…]
@@
If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$
IV Tree-based RL training
Leveraging the semantic features extracted from the raw visual inputs courtesy of the perception stack, we now focus on training an interpretable RL model that maps semantic features to action outputs, as shown in Fig. 2. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step.
Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [5] as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization. Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions. The detailed RL training is listed in Algorithm 3
IV-A Interpretable control tree